Use mathimatical Induction to prove that
1²+2²+3²+.......n²= n(n+1)(2n+1) divided by 6
1
Expert's answer
2020-12-01T02:11:12-0500

First we need to show the result is true for smallest value of n, in this case n = 0/

When n =0 the left side has only one term n2 = 0

The right side is

"\frac{n(n+2)(2n+1)}{6} = \frac{0(0+1)(0+1)}{6} = 0"

The second step is to show the statement is true for n = k, k+1 and so on.

Suppose, that statement is true for n = K.

12 + 22 + 32 + … + k2 = "\frac{k(k+1)(2k+1)}{6}"

We need to prove, that statement is true fof n = k+1

12 + 22 + 32 + … + (k+1)2 = "\frac{(k+1)(k+2)(2(k+1) +1)}{6}"

12 + 22 + 32 + … +k2 + (k+1)2 = "\frac{(k+1)(k+2)(2(k+1) +1)}{6}"

We know, that

12 + 22 + 32 + … +k2 = "\frac{k(k+1)(2k+1)}{6}"

So,

"\frac{k(k+1)(2k+1)}{6} + (k+1)^2 = \frac{k(k+1)(2k+1) +6(k+1)^2}{6}"

"= \frac{k+1}{6}(k(2k+1) + 6(k+1))"

"= \frac{k+1}{6}(2k^2 + k + 6k + 6)"

"= \frac{k+1}{6}(2k^2 + 7k + 6)"


"2k^2 + 7k + 6 = 0"

"2k^2 + 4k + 3k + 6 = 0"

"2k(k+2) + 3(k+2) =0"

"(2k+3)(k+2)=0"


"frac{k+1}{6}(2k^2 + 7k + 6) = \frac{(k+1)(k+2)(2k+3)}{6}"

"= \frac{(k+1)((k+1)+1)(2(k+1)+1)}{6}" = RHS hence proved